# KernelDNA: Dynamic Kernel Sharing via Decoupled Naive Adapters

### v1

@@ 6.1 Why Kernel Sharing is Effective?
While experiments validate the effectiveness of KernelDNA across multiple vision tasks, further exploration into its underlying mechanisms is critical. We conduct Linear CKA analysis on ResNet18 trained with KernelDNA under optimal configurations on ImageNet-1K (Fig. 6). The left subplot shows the similarity grid of the original ResNet18 layers, while the middle subplot represents KernelDNA’s results, where parameter-sharing replaces or extends original convolutional layers, resulting in slightly more layers. The right subplot compares layer-wise similarity between KernelDNA and the original model. We observe that:
@@
Block-Wise Specificity: KernelDNA retains intra-block commonalities (shared patterns) between parent and child layers while introducing inter-block heterogeneity (distinct adaptations), aligning with principles of epigenetic regulation in feature learning.
@@ 6.2 Do Adapters Really Learn the Differences?
To further validate whether our proposed “DNA” adapter—equipped with attention mechanisms—genuinely learns diversity across shared child layers, we visualize the spatial attention features, as shown in Fig. 7. The figure compares three child layers derived from the third parent layer and three from the fourth parent layer. The results demonstrate that child layers from the same parent exhibit distinct variations, confirming that our adapter preserves shared commonality features from parent kernels while introducing differentiated diversity among child layers. This aligns with our hypothesis that the adapter balances inheritance and specialization, ensuring both parameter efficiency and expressive power.
7 Limitations
While our proposed kernel-sharing strategy achieves significant parameter reduction, it does not reduce computational costs (i.e., FLOPs) and may introduce marginal additional computational. Though this trade-off might make it less ideal for compute-constrained devices, our method remains broadly applicable given that modern hardware platforms are predominantly memory-bound rather than computation-limited. By prioritizing parameter efficiency without compromising the inference speed of standard convolutions, KernelDNA aligns with the practical demands of most real-world deployment scenarios.
@@ 8 Conclusion
KernelDNA introduces a lightweight convolution plugin that addresses the parameter and efficiency limitations of dynamic convolutions. By leveraging cross-layer weight sharing and adapter-based modulation, it combines dynamic kernel adaptation with standard convolution’s efficiency, reducing parameters while enhancing performance. Experiments show state-of-the-art accuracy-efficiency trade-offs on image classification and dense prediction tasks, outperforming existing methods. KernelDNA maintains hardware-friendly inference speeds (e.g., 90–99% of base models’ throughput) with 1.2–5 fewer parameters than SOTA variants, offering a practical solution for adaptive CNNs without accuracy compromises.

### v2

@@ 6.1 Why Kernel Sharing is Effective?
While experiments validate the effectiveness of KernelDNA across multiple vision tasks, further exploration into its underlying mechanisms is critical. We conduct Linear CKA analysis on ResNet18 trained with KernelDNA under optimal configurations on ImageNet-1K (Figure 5). The left subplot shows the similarity grid of the original ResNet18 layers, while the middle subplot represents KernelDNA’s results, where parameter-sharing replaces or extends original convolutional layers, resulting in slightly more layers. The right subplot compares layer-wise similarity between KernelDNA and the original model. We observe that:
@@
Block-Wise Specificity: KernelDNA retains intra-block commonalities (same parent) between parent and child layers while introducing inter-block heterogeneity (different parents), aligning with principles of epigenetic regulation in feature learning.
@@ 6.2 Do Adapters Really Learn the Differences?
To further validate whether our proposed “DNA” adapter—equipped with attention mechanisms—genuinely learns diversity across shared child layers, we visualize the spatial attention features, as shown in Figure 6. The figure compares three child layers derived from the third parent layer and three from the fourth parent layer. The results demonstrate that child layers from the same parent exhibit distinct variations, confirming that our adapter preserves shared commonality features from parent kernels while introducing differentiated diversity among child layers. This aligns with our hypothesis that the adapter balances inheritance and specialization, ensuring both parameter efficiency and expressive power.
@@ 7 Conclusion
KernelDNA introduces a lightweight convolution plugin that addresses the parameter and efficiency limitations of dynamic convolutions. By leveraging cross-layer weight sharing and adapter-based modulation, it combines dynamic kernel adaptation with standard convolution’s efficiency, reducing parameters while enhancing performance. Experiments show state-of-the-art (SOTA) accuracy-efficiency trade-offs on image classification and dense prediction tasks, outperforming existing methods. KernelDNA maintains hardware-friendly inference speeds with 1.2–5 fewer parameters than previous SOTA variants, offering a practical solution for adaptive CNNs with accuracy profitability.
Supplementary Material
A Limitations
Despite the substantial parameter reduction achieved by our kernel-sharing strategy, there are some limitations to consider. KernelDNA may introduce a slight increase in computational overhead (e.g., marginally higher FLOPs) due to the adapter modules, although our experiments demonstrate that the impact on inference speed is minimal—maintaining 90–99% throughput of the base models. This trade-off may be less suitable for extremely compute-constrained environments. However, given that most modern hardware is memory-bound rather than compute-bound, the parameter efficiency and hardware-friendly design of KernelDNA makes it well-suited for a wide range of real-world applications.
B Additional Experiments
B.1 Channel Reduction and Expansion
To further investigate the flexibility of KDNA, we conduct experiments with various channel reduction and expansion ratios, as summarized in Table 7. The results demonstrate that the default configuration (highlighted in bold) achieves the optimal balance between parameter efficiency and inference speed, with only marginal accuracy changes observed as the ratio varies.
| Type | Ratio | Params | TP GPU (fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| Channel | 1 | 10.719M | 7670 | 74.28 | 92.42 |
| 1/2 | 9.733M | 7678 | 74.25 | 92.41 | |
| 1/4 | 9.241M | 7683 | 74.23 | 92.38 | |
| 1/8 | 8.994M | 7688 | 74.16 | 92.26 | |
| 1/16 | 8.871M | 7691 | 73.76 | 91.91 | |
| Spatial | 1 | 9.241M | 7683 | 74.23 | 92.38 |
| 2 | 9.241M | 7682 | 74.22 | 92.36 | |
| 4 | 9.241M | 7678 | 74.19 | 92.29 | |
| 8 | 9.241M | 7670 | 74.21 | 92.30 | |
| 16 | 9.242M | 7665 | 74.11 | 92.23 | |
| Filter | 1 | 9.241M | 7683 | 74.23 | 92.38 |
| 2 | 9.243M | 7679 | 74.21 | 92.35 | |
| 4 | 9.248M | 7675 | 74.18 | 92.29 | |
| 8 | 9.257M | 7671 | 74.19 | 92.31 | |
| 16 | 9.275M | 7662 | 74.12 | 92.26 |